1. Providing for Doubtful Accounts At the end of the current year, the accounts receivable account...

1.

Providing for Doubtful Accounts

At the end of the current year, the accounts receivable account has a debit balance of $755,000 and sales for the year total $8,560,000.

  1. The allowance account before adjustment has a debit balance of $10,200. Bad debt expense is estimated at 3/4 of 1% of sales.
  2. The allowance account before adjustment has a debit balance of $10,200. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $32,600.
  3. The allowance account before adjustment has a credit balance of $7,000. Bad debt expense is estimated at 1/2 of 1% of sales.
  4. The allowance account before adjustment has a credit balance of $7,000. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $58,100.

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.

a. $
b. $
c. $
d.

$

2.

Determine the amount to be paid in full settlement of each invoice, assuming that credit for returns and allowances was received prior to payment and that all invoices were paid within the discount period.

Merchandise Freight Paid
by Seller
Freight Terms Returns and
Allowances
a. $9,100      $318      FOB Shipping Point, 1/10, net 30 $900     
b. $10,300      $61      FOB Destination, 2/10, net 45 $2,300     
a. $
b. $

3.

Using the allowance method of accounting for uncollectible receivables.

Transactions:
April 1 Sold merchandise on account to Jim Dobbs, $8,900. The cost of the merchandise is $3,560.
June 10 Received payment for one-third of the receivable from Jim Dobbs and wrote off the remainder.
Oct. 11 Reinstated the account of Jim Dobbs and received cash in full payment.

Required:

Journalize the above transactions. Refer to the Chart of Accounts for exact wording of account titles. Round your answers to nearest dollar amount.
Answer #1

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.

a. $8560000*1%*3/4 = 64200
b. $32600+10200 = 42800
c. $8560000*1%*1/2 = 42800
d.

$58100-7000 = 51100

2a) Amount paid =  (9100-900)*99%+318 = 8436

Amount paid = (10300-2300)*98% = 7840

3) Journal entry

Date account and explanation Debit Credit
Apr 1 Account receivable 8900
Sales revenue 8900
Cost of goods sold 3560
Inventory 3560
June 10 Cash (8900/3) 2967
Allowance for doubtful account 5933
Account receivable 8900
oct 11 Account receivable 5933
Allowance for doubtful account 5933
Cash 5933
Account receivable 5933